OpenStudy (anonymous):

The cost of printing a flyer to spread knowledge of an organization that you run is a flat rate of $9 plus $0.05 per copy. Let C represent the total cost of printing and x represent the number of copies. Write in symbolic form an equation that will relate C and x. Organize the data into a table of values. Begin with 500 copies, increase by increments of 500, and end with 2500 copies. How much would it cost to print 2500 flyers and what equation would you use?a. C = 9 + 0.05x 2500 flyers will cost $159c C = 9.05x 2500 flyers will cost $27,150b C = 9 + 0.05x 2500 flyers will cost $134

OpenStudy (texaschic101):

C = 9 + .05x <-- this is the equation you use to find the cost for 500 copies.. for 2000 copies C = 9 + .05(500) C = 9 + .05(2000) C = 9 + 25 C = 9 + 100 C = 34 C = 109 for 1000 copies for 2500 copies C = 9 + .05(1000) C = 9 + .05(2500) C = 9 + 50 C = 9 + 125 C = 59 C = 134 for 1500 copies C = 9 + .05(1500) C = 9 + 75 C = 84 TABLE --------------------------------------- x -- 500 -- 1000 -- 1500 -- 2000 -- 2500 ---------------------------------------- C -- 34 ---- 59 -----84-----109----134 ----------------------------------------